OpenStudy (anonymous):

2+3=8 3+7=27, 4+5=32, 5+8=60, 6+7=72, 7+8=?? Solve it ?

OpenStudy (anonymous):

MUAHAHAHA

OpenStudy (anonymous):

haha hard one xD

OpenStudy (anonymous):

2+3 = 5 5*2-2 = 8 3+7 = 10 10*3 = 30-3 = 27 Do you see the pattern here?

OpenStudy (anonymous):

lol i know the answer

OpenStudy (anonymous):

Oh did you understand it though?

OpenStudy (anonymous):

It should be 98